40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3966     Entry time: Mon Nov 22 18:39:53 2010
Author: Jenne 
Type: Update 
Category: SUS 
Subject: ETMU07: magnets glued to optic. ETMU05: magnets removed 

[Suresh, Jenne]

A story about minor disasters, and crises averted:

Once upon a time, in a cleanroom not so far away..... there lived an optic.  To preserve anonymity, we shall call him "ETMU05".  This optic had a rough day.  When removing the grippers from the magnet-to-optic fixture, 4 out of 6 magnets broke off the dumbbells (the dumbbells were still securely glued to the optic...these had come out of the same batch that had problems last week, same problem).  The remaining 2, LL and LR, were sadly of the same polarity.  This is bad, because it means that the "humans" taking care of "ETMU05" didn't check the polarity of the face magnets properly, and ensure that they were laid out in an every-other pattern (LL and UR having the same polarity, and LR and UL having the opposite).  So, the humans removed all magnets and dumbbells from ETMU05.  All remaining glue was carefully scrubbed off the surfaces of ETMU05 using lens paper and acetone, and the magnets and dumbbells were sonicated in acetone, scrubbed with a lint-free wipe, sonicated again, and then scrubbed again to remove the glue.  ETMU05 had a nice cleansing, and was drag wiped on both the AR and HR surfaces with acetone and iso.  ETMU05 is now on vacation in a nice little foil hut.

His friend, (let's call him ETMU07) had a set of magnets (with polarities carefully confirmed) glued to him.  The cleaned magnets and dumbbells removed from ETMU05 were reglued to their dumbbells, and should be dry by tomorrow. 

.....And then they lived happily ever after.  The End.

 


The revised schedule / status table:

StatusTable.png

ELOG V3.1.3-